In 1987, a gargantuan iceherg broke away from the Ross Ioe Sheet in Antarctica It was approximadly a rectangle wih dimensioes 154 km long 35.0 km widk, and 230.0 m thick What is the mass of this iceherg, given that the demity of ice is 917 kg/m kg How much heal trandfer in joukcs is needed to mell the iceherg? How many years wowuld i take sunlight alone to metl sce this thick i the ice abts an avcrage of 109 W/m2. 12.0h per day?
media%2F3bc%2F3bce7c34-decc-48d7-9d65-51
media%2Fdcd%2Fdcd520b4-4523-44a9-be5e-aa
media%2Fd8f%2Fd8f87c3d-d305-4d5d-abb8-ce
Answer #1

1) the heat capacity is the amount of heat absrobed by a substance in increaseing its temperature by 10C

Given: Heat absrobed = 3851 Joules

The increase in temperature = 43.3 0C

Heat capacity = Heat absorbed / increase in temperature = 3851 / 43.3 = 88.94 J / 0C

Specific heat of a substance= heat required to raise the temprature of one gram of substance by 10C

Given mass =371 g

So specific heat capacity = 88.94 J / 0C / 371g = 0.240J /g 0C

So the object is compsed of silver

2) Overall balanced equation will be

Cu(s) + 2Ag+(aq) ---> Cu+2(aq) + 2Ag(s)

3) The heat provided to ice first will be used for raising the temperature, if further heat is left then it will be used for conversion of ice to water (l) and rest for heating water

Heat given = 687 cal = 687 X 4.18 J = 2876.33 Joules

Heat given = heat absorbed = Mass of ice X specific heat of ice X change in temperature

Heat absorbed by ice = Q1 = 5 X 2.087 J/g 0C X change in temperature = 5 X 2.087 X (20-0) = 208.7 Joules

Heat absorbed in conversion of ice to water = Mass X Heating capacity = 5 X 333.6 J /g = 1668 Joules

Heat left = 2876.33 - (208.7 + 1668) = 999.63 Joules

This heat will be absrobed for raising the temperature of water

Heat = Mass of water X specific heat of water X change in temperature

999.63 = 5 X 4.184 X Change in temperature

Change in temperature = 47.78 0C

Final temperature of water = 47.78 0C